# Efficient GW band structure calculations using Gaussian basis functions and application to atomically thin transition-metal dichalcogenides

### v2

@@ V.5 Reference GW calculations with the projector-augmented-wave scheme (VASP)
The average deviations of $GW$ band gaps between the numerical approaches are summarized in Fig. 5. The discrepancies may be attributed to several factors, including the use of different pseudopotentials, the limited size of the aug-DZVP-MOLOPT basis compared to high plane-wave cutoffs, and sensitivity to convergence parameters in frequency integration or dielectric matrix evaluation. Time-frequency resolution may also contribute to residual differences, which could be mitigated using Hedin’s shift [112]. This is also demonstrated in Appendix J, where we report some available results in the literature for the band gaps of the TMDCs at the $G_{0}W_{0}$@PBE+SOC. As we have stated earlier, one should note however that the comparison with the already existing calculations is fairly difficult as most of these use different lattice constants, which has a huge impact on the band gap [80].
Comparison to experimental measurements is inherently challenging, as the band gap is highly sensitive to external influences that are difficult to control, such as substrate screening effects [113] and strain [80]. Reported experimental band gaps for the four materials range from 1.9 eV to 2.4 eV [114, 113, 115, 116], which aligns with our $GW$ results (Table 1). However, achieving a precise one-to-one correspondence for each material remains elusive at this stage, especially given that many experimentally reported band gaps are measured in the presence of a substrate, which is not considered in this work, as it lies outside the scope of this manuscript.
For completeness, we also performed benchmark full-spinor calculations with BerkeleyGW using the generalized plasmon pole model [97], see Table 2 in Appendix I. This approximation yields bandgaps with significant deviations - up to $0.25$ eV - when compared with the full frequency calculation of the quasi-particle self-energy. These discrepancies sharply contrast with the excellent agreement of below $50$ meV between the different codes, as reported in Fig. 5.
@@
Figure 6: Scaling of the computational time for the $G_{0}W_{0}$@PBE+SOC band gap calculation of monolayer MoS2 with respect to the number of cores. We use an aug-SZV-MOLOPT basis and light settings (Sec. V.3).
@@ IX Conclusion and outlook
We have developed a $GW$ space-time algorithm for periodic systems using Gaussian basis functions with spin-orbit coupling, enabling accurate quasiparticle band structure calculations for atomically thin materials. Our implementation, available in the open-source code CP2K, achieves high accuracy: for monolayer MoS2, MoSe2, WS2, and WSe2, $GW$ band gaps agree on average within 50 meV to plane-wave $GW$ band gaps from BerkeleyGW and VASP. Using lighter computational settings, the average agreement is 70 meV, and a complete $GW$ band structure calculation can be performed on a laptop (Intel i5, 192GB RAM) in approximately 24 hours, or in less than 30 minutes on 1024 HPC cores. Future extensions of this work will combine the present $GW$ algorithm with real-space grids for $G$ and $W$ [117, 52, 118]. We expect this to significantly reduce the number of required lattice summations, further lowering the computational cost. The Gaussian basis sets investigated in this work yield converged band gaps and band structures within $\sim$ 100 meV. Developing Gaussian basis sets that enable convergence of periodic $GW$ calculations ideally within 10 meV is subject of ongoing work.
@@
Inputs and outputs of all calculations reported in this work are available in a Github repository [119]. The algorithm presented in this work is available in the open-source package CP2K [84, 82, 83].
@@ Appendix A Periodic RI
For computing $\chi_{PQ}^{\mathbf{R}}(i\tau)$ from Eq. (42), the three-centre integrals $(\mu\mathbf{R}\,\nu\mathbf{T}|P\mathbf{R})$ are appearing. Moreover, in Eq. (41) the metric matrix $M_{PQ}(\mathbf{k})$ is appearing. $(\mu\mathbf{R}\,\nu\mathbf{T}|P\mathbf{R})$ and $M_{PQ}(\mathbf{k})$ are stemming from the application of the resolution of the identity (RI) [66, 44, 47]. In this Appendix, we give details on RI with periodic boundary conditions.
@@
where the product of two Gaussian functions $\phi_{\mu}(\mathbf{r})$ and $\phi_{\nu}(\mathbf{r})$ is approximated as a linear combination of auxiliary Gaussian functions $\varphi_{P}(\mathbf{r})$ with coefficients $B_{P}^{\mu\nu}$. $B_{P}^{\mu\nu}$ are determined by fitting procedures like least-squares minimization [66] to get a good approximation in Eq. (65).
@@
We show in this Appendix that the expansion coefficients $B_{P\mathbf{P}}^{\mu\mathbf{R}\nu\mathbf{T}}$ are determined by least-squares minimization under a metric $m(\mathbf{r})$ [66] as
@@
In the following, we outline our proof of Eq. (67). As in non-periodic RI [66], we define the residual $\mathfrak{R}$ of Eq. (66),
@@
For deriving the irreducible response $\chi^{\mathbf{R}}_{PQ}(i\tau)$ for a periodic system, Eq. (42), we start from Eq. (3) and (2) for computing $\chi$ for a molecule. For a molecule, we sum over the molecular quantum number $i$ and $a$ for occupied and empty MOs in Eq. (2). For a periodic system, the quantum numbers of the one-electronic states are $i\mathbf{k}$ and $a\mathbf{k}$ labeling an occupied and empty band at $k$-point $\mathbf{k}$ in the BZ. We thus replace and abbreviate
@@
and $\delta$ is the filter threshold. All other quantities of the algorithm from Fig. 7 are defined in the main text. The three tensor operations are executed by the sparse-tensor library dbt [120].
@@ Appendix D Lattice sum of the Coulomb matrix element (46)
by explicit summation of $\sum_{\mathbf{R}}$. Several schemes for executing this lattice sum have been described, see Ref. [121], Appendix of Ref. [73] and references therein.
We first note that the lattice summation in Eq. (90) is divergent if $\varphi_{P}^{\mathbf{0}}$ and $\varphi_{Q}^{\mathbf{R}}$ are $s$-type basis functions in the limit $\mathbf{k}\hskip 1.00006pt{\rightarrow}\hskip 1.00006pt0$. This is because $(\varphi_{P}^{\mathbf{0}}|\varphi_{Q}^{\mathbf{R}})\hskip 1.69998pt{\sim}\hskip 1.69998pt|\mathbf{R}|^{-1}$ and the lattice sum $\sum\limits_{\mathbf{R}}(\varphi_{P}^{\mathbf{0}}|\varphi_{Q}^{\mathbf{R}})\hskip 1.69998pt{\sim}\hskip 1.69998pt\sum\limits_{\mathbf{R}}|\mathbf{R}|^{-1}$ (for $\mathbf{k}\hskip 1.69998pt{=}\hskip 1.69998pt0$) can be approximated as an integral $\int_{\mathbb{R}^{d}}d\mathbf{R}\,|\mathbf{R}|^{-1}\hskip 1.69998pt{\sim}\hskip 1.69998pt\int_{0}^{\infty}dR\,R^{d-1}\,R^{-1}$ which diverges for dimensionality $d\hskip 1.69998pt{=}\hskip 1.69998pt1,2,3$ (1d-molecular chains, 2d-materials or surfaces, 3d-bulk solids). More precisely, the Coulomb matrix element $V_{PQ}(\mathbf{k})$ diverges at $\mathbf{k}\hskip 1.00006pt{\rightarrow}\hskip 1.00006pt0$ as $V_{PQ}(\mathbf{k})\hskip 1.69998pt{\sim}\hskip 1.69998pt|\mathbf{k}|^{-(d-1)}$ for $s$-functions $P,Q$ for $d\hskip 1.69998pt{=}\hskip 1.69998pt2,3$ (for $d\hskip 1.69998pt{=}\hskip 1.69998pt1$ the divergence is logarithmic). [76] The whole $GW$ algorithm is not divergent because this $1/k^{d-1}$ divergence is integrable in the Brillouin zone, see as an example Eq. (48): $\int_{\text{BZ}}d\mathbf{k}\,1/k^{d-1}\hskip 1.69998pt{\sim}\hskip 1.69998pt\int_{0}^{k_{\text{max}}}dk\,k^{d-1}/k^{d-1}$, where $k_{\text{max}}$ is determined by the Brillouin zone size.
The requirements for convergence of the lattice sum (90) have been extensively discussed in the literature [121, 73] and references therein. We reproduce the arguments here to make the discussion self-contained. We consider the lattice sum
@@
where $2^{k_{\text{min}}}$ and $2^{l_{\text{min}}}$ are the lowest nonvanishing dipole moments of $\varphi^{\mathbf{0}}_{P}$ and $\rho^{\mathbf{T}}$, respectively [121]. In our case, $\varphi^{\mathbf{0}}_{P}$ can be an $s$-type basis function, i.e., $k_{\text{min}}\hskip 1.69998pt{=}\hskip 1.69998pt0$, so
@@
We start the derivation of an absolutely convergent lattice summation by defining a $(2N_{1})\hskip 0.89996pt{\times}\hskip 0.59998pt(2N_{2})$ supercell "SC" as sketched in Fig. 8. Here, $N_{1}$ and $N_{2}$ correspond to the $N_{1}\hskip 0.89996pt{\times}\hskip 0.59998ptN_{2}$ Monkhorst-Pack $k$-point mesh [65] used for the BZ integration (48) of $W$. A lattice vector $\mathbf{R}$ belongs to SC if
@@ Appendix F Proof that absolutely convergent lattice sum (109) reproduces plane-wave GW Coulomb integrals
For 2D systems, the Coulomb operator is truncated in the non-periodic direction to eliminate artificial image interactions [100] and all following steps are analogous.
@@
Note that Eq. (118) has been used in Ref. [122] to compute the Coulomb matrix element $V_{PQ}(\mathbf{k})$ of Gaussians $P,Q$.
@@ Appendix H Spin-orbit coupling from HGH pseudopotentials
We employ spin-orbit coupling (SOC) from Hartwigsen-Goedecker-Hutter (HGH) pseudopotentials [54, 56],
@@
To demonstrate the effects of this correction, we provide in the following the PBE band structures of all monolayers with and without SOC (Fig. 10). We observe that SOC splits the top valence bands, predominantly at the K-point, in agreement with the literature [9].
@@ Appendix I GW band gaps with plasmon-pole model (BerkeleyGW)
In this Appendix, we present additional data of the band gaps for the TMDCs benchmarked in this work using BerkeleyGW (Table 2). We compare the generalized plasmon pole model (GPP) [97] with the full frequency implentation (data of the full-frequency calculation is reproduced from Table 1). Some available results in the literature are given in Table 4.
@@
C2DB [123, 124] GPAW PAW 18.1 3.184 12x12x1 N/A 59 $\infty$111extrapolated from 11, 13 and 15 Ry 2.53
Pela et al. [125] exciting LAPW+lo 16.0 3.190 18x18x1 400 N/A N/A 2.45
Echeverry et al. [126] VASP PAW 17.0 3.220 12x12x1 600 N/A N/A 2.31
Qiu et al. [127] Berkeley GW NC 25.0 3.150 24x24x1 6000 350 35 2.45
@@
This work CP2K (TZV2P) NC 15.0 3.320 32x32x1 1.99
This work CP2K (aug-DZVP) NC 15.0 3.320 32x32x1 1.94
@@
C2DB [123, 124] GPAW PAW 18.3 3.320 9x9x1 N/A 59 $\infty$1 2.12
Echeverry et al. [126] VASP PAW 17.0 3.320 12x12x1 600 N/A N/A 2.13
@@
This work CP2K (TZV2P) NC 15.0 3.186 32x32x1 2.44
This work CP2K (aug-DZVP) NC 15.0 3.186 32x32x1 2.34
@@
C2DB [123, 124] GPAW PAW 18.1 3.186 12x12x1 N/A 59 $\infty$1 2.53
@@
This work CP2K (TZV2P) NC 15.0 3.319 32x32x1 2.02
This work CP2K (aug-DZVP) NC 15.0 3.319 32x32x1 1.93
@@
C2DB [123, 124] GPAW PAW 18.3 3.319 12x12x1 N/A 59 $\infty$1 2.13
Echeverry et al. [126] VASP PAW 17.0 3.280 12x12x1 600 N/A N/A 2.06
@@
Gillen et al. [128] Berkeley GW NC 25.0 3.150 24x24x1 200 88 22 2.69
Kim et al. [129] Berkeley GW NC 25.0 3.190 12x12x1 3000 125 35 2.40
Qiu et al. [127] Berkeley GW NC 25.0 3.150 24x24x1 6000 350 35 2.59
@@
Gillen et al. [128] Berkeley GW NC 25.0 3.279 24x24x1 200 88 22 2.33
Kim et al. [129] Berkeley GW NC 25.0 3.316 12x12x1 3000 125 35 2.08
@@
Figure 11: $G_{0}W_{0}$@PBE+SOC band gap of monolayer MoS2 and execution time as a function of the number of time points $\tau$ (Sec. IV), the $k$-mesh (Eqs. (21), (23), (32)), the filter threshold (for Eqs. (42), (56) and (59), see Eqs. (88) and (89)), the simulation cell box height (Sec. V.3), the size factor $\Delta$ for $V_{PQ}(\mathbf{k})$ (Eq. (110)) and the number of basis functions (Sec. V.3). Default parameters are reported on top. In (a), we also show $G_{0}W_{0}$@PBE+SOC with Hedin’s shift [1, 107, 108, 109, 3, 110] to avoid poles of the self-energy close to the quasiparticle solution [111, 112].

### v3

@@ V.5 Reference GW calculations with the projector-augmented-wave scheme (VASP)
The average deviations of $GW$ band gaps between the numerical approaches are summarized in Fig. 5. The discrepancies may be attributed to several factors, including the use of different pseudopotentials, the limited size of the aug-DZVP-MOLOPT basis compared to high plane-wave cutoffs, and sensitivity to convergence parameters in frequency integration or dielectric matrix evaluation. Time-frequency resolution may also contribute to residual differences, which could be mitigated using Hedin’s shift [Schambeck2024]. This is also demonstrated in Appendix J, where we report some available results in the literature for the band gaps of the TMDCs at the $G_{0}W_{0}$@PBE+SOC. As we have stated earlier, one should note however that the comparison with the already existing calculations is fairly difficult as most of these use different lattice constants, which has a huge impact on the band gap [Zollner2019].
Comparison to experimental measurements is inherently challenging, as the band gap is highly sensitive to external influences that are difficult to control, such as substrate screening effects [Zhang2016] and strain [Zollner2019]. Reported experimental band gaps for the four materials range from 1.9 eV to 2.4 eV [Ryou2016, Zhang2016, Hill2016, Yankowitz2015], which aligns with our $GW$ results (Table 1). However, achieving a precise one-to-one correspondence for each material remains elusive at this stage, especially given that many experimentally reported band gaps are measured in the presence of a substrate, which is not considered in this work, as it lies outside the scope of this manuscript.
For completeness, we also performed benchmark full-spinor calculations with BerkeleyGW using the generalized plasmon pole model [Hybertsen1986], see Table 2 in Appendix I. This approximation yields bandgaps with significant deviations - up to $0.25$ eV - when compared with the full frequency calculation of the quasi-particle self-energy. These discrepancies sharply contrast with the excellent agreement of below $50$ meV between the different codes, as reported in Fig. 5.
@@
Figure 6: Scaling of the computation time for the $G_{0}W_{0}$@PBE+SOC band gap calculation of monolayer MoS2 with respect to the number of cores. We use an aug-SZV-MOLOPT basis and light settings (Sec. V.3).
@@ IX Conclusion and outlook
We have developed a $GW$ space-time algorithm for periodic systems using Gaussian basis functions with spin-orbit coupling, enabling accurate quasiparticle band structure calculations for atomically thin materials. Our implementation, available in the open-source code CP2K, achieves high accuracy: for monolayer MoS2, MoSe2, WS2, and WSe2, $GW$ band gaps agree on average within 50 meV to plane-wave $GW$ band gaps from BerkeleyGW and VASP. Using lighter computational settings, the average agreement is 70 meV, and a complete $GW$ band structure calculation can be performed on a laptop (Intel i5, 192GB RAM) in approximately 24 hours, or in less than 30 minutes on 1024 HPC cores. Future extensions of this work will combine the present $GW$ algorithm with real-space grids for $G$ and $W$ [duchemin2019, duchemin2021, Delesma2024]. We expect this to significantly reduce the number of required lattice summations, further lowering the computational cost. The Gaussian basis sets investigated in this work yield converged band gaps and band structures within $\sim$ 100 meV. Developing Gaussian basis sets that enable convergence of periodic $GW$ calculations ideally within 10 meV is subject of ongoing work.
@@
Inputs and outputs of all calculations reported in this work are available in a Github repository [github_repo]. The algorithm presented in this work is available in the open-source package CP2K [cp2k, Kuehne2020, Kuehne2025].
@@ Appendix A Periodic RI
For computing $\chi_{PQ}^{\mathbf{R}}(i\tau)$ from Eq. (42), the three-centre integrals $(\mu\mathbf{R}\,\nu\mathbf{T}|P\mathbf{R})$ are appearing. Moreover, in Eq. (41) the metric matrix $M_{PQ}(\mathbf{k})$ is appearing. $(\mu\mathbf{R}\,\nu\mathbf{T}|P\mathbf{R})$ and $M_{PQ}(\mathbf{k})$ are stemming from the application of the resolution of the identity (RI) [Vahtras1993, zhu2021all, Graml2024]. In this Appendix, we give details on RI with periodic boundary conditions.
@@
where the product of two Gaussian functions $\phi_{\mu}(\mathbf{r})$ and $\phi_{\nu}(\mathbf{r})$ is approximated as a linear combination of auxiliary Gaussian functions $\varphi_{P}(\mathbf{r})$ with coefficients $B_{P}^{\mu\nu}$. $B_{P}^{\mu\nu}$ are determined by fitting procedures like least-squares minimization [Vahtras1993] to get a good approximation in Eq. (65).
@@
We show in this Appendix that the expansion coefficients $B_{P\mathbf{P}}^{\mu\mathbf{R}\nu\mathbf{T}}$ are determined by least-squares minimization under a metric $m(\mathbf{r})$ [Vahtras1993] as
@@
In the following, we outline our proof of Eq. (67). As in non-periodic RI [Vahtras1993], we define the residual $\mathfrak{R}$ of Eq. (66),
@@
For deriving the irreducible response $\chi^{\mathbf{R}}_{PQ}(i\tau)$ for a periodic system, Eq. (42), we start from Eq. (3) and (2) for computing $\chi$ for a molecule. For a molecule, we sum over the molecular quantum numbers $i$ and $a$ for occupied and empty MOs in Eq. (2). For a periodic system, the quantum numbers of the one-electronic states are $i\mathbf{k}$ and $a\mathbf{k}$ labeling an occupied and empty band at $k$-point $\mathbf{k}$ in the BZ. We thus replace and abbreviate
@@
and $\delta$ is the filter threshold. All other quantities of the algorithm from Fig. 7 are defined in the main text. The three tensor operations are executed by the sparse-tensor library dbt [Seewald2020].
@@ Appendix D Lattice sum of the Coulomb matrix element (46)
by explicit summation of $\sum_{\mathbf{R}}$. Several schemes for executing this lattice sum have been described, see Ref. [Piela1982], Appendix of Ref. [Grundei2017] and references therein.
We first note that the lattice summation in Eq. (90) is divergent if $\varphi_{P}^{\mathbf{0}}$ and $\varphi_{Q}^{\mathbf{R}}$ are $s$-type basis functions in the limit $\mathbf{k}\hskip 1.00006pt{\rightarrow}\hskip 1.00006pt0$. This is because $(\varphi_{P}^{\mathbf{0}}|\varphi_{Q}^{\mathbf{R}})\hskip 1.69998pt{\sim}\hskip 1.69998pt|\mathbf{R}|^{-1}$ and the lattice sum $\sum\limits_{\mathbf{R}}(\varphi_{P}^{\mathbf{0}}|\varphi_{Q}^{\mathbf{R}})\hskip 1.69998pt{\sim}\hskip 1.69998pt\sum\limits_{\mathbf{R}}|\mathbf{R}|^{-1}$ (for $\mathbf{k}\hskip 1.69998pt{=}\hskip 1.69998pt0$) can be approximated as an integral $\int_{\mathbb{R}^{d}}d\mathbf{R}\,|\mathbf{R}|^{-1}\hskip 1.69998pt{\sim}\hskip 1.69998pt\int_{0}^{\infty}dR\,R^{d-1}\,R^{-1}$ which diverges for dimensionality $d\hskip 1.69998pt{=}\hskip 1.69998pt1,2,3$ (1d-molecular chains, 2d-materials or surfaces, 3d-bulk solids). More precisely, the Coulomb matrix element $V_{PQ}(\mathbf{k})$ diverges at $\mathbf{k}\hskip 1.00006pt{\rightarrow}\hskip 1.00006pt0$ as $V_{PQ}(\mathbf{k})\hskip 1.69998pt{\sim}\hskip 1.69998pt|\mathbf{k}|^{-(d-1)}$ for $s$-functions $P,Q$ for $d\hskip 1.69998pt{=}\hskip 1.69998pt2,3$ (for $d\hskip 1.69998pt{=}\hskip 1.69998pt1$ the divergence is logarithmic). [FREYSOLDT20071] The whole $GW$ algorithm is not divergent because this $1/k^{d-1}$ divergence is integrable in the Brillouin zone, see as an example Eq. (48): $\int_{\text{BZ}}d\mathbf{k}\,1/k^{d-1}\hskip 1.69998pt{\sim}\hskip 1.69998pt\int_{0}^{k_{\text{max}}}dk\,k^{d-1}/k^{d-1}$, where $k_{\text{max}}$ is determined by the Brillouin zone size.
The requirements for convergence of the lattice sum (90) have been extensively discussed in the literature [Piela1982, Grundei2017] and references therein. We reproduce the arguments here to make the discussion self-contained. We consider the lattice sum
@@
where $2^{k_{\text{min}}}$ and $2^{l_{\text{min}}}$ are the lowest nonvanishing dipole moments of $\varphi^{\mathbf{0}}_{P}$ and $\rho^{\mathbf{T}}$, respectively [Piela1982]. In our case, $\varphi^{\mathbf{0}}_{P}$ can be an $s$-type basis function, i.e., $k_{\text{min}}\hskip 1.69998pt{=}\hskip 1.69998pt0$, so
@@
We start the derivation of an absolutely convergent lattice summation by defining a $(2N_{1})\hskip 0.89996pt{\times}\hskip 0.59998pt(2N_{2})$ supercell "SC" as sketched in Fig. 8. Here, $N_{1}$ and $N_{2}$ correspond to the $N_{1}\hskip 0.89996pt{\times}\hskip 0.59998ptN_{2}$ Monkhorst-Pack $k$-point mesh [Monkhorst1976] used for the BZ integration (48) of $W$. A lattice vector $\mathbf{R}$ belongs to SC if
@@ Appendix F Proof that absolutely convergent lattice sum (109) reproduces plane-wave GW Coulomb integrals
For 2D systems, the Coulomb operator is truncated in the non-periodic direction to eliminate artificial image interactions [Ismail2006] and all following steps are analogous.
@@
Note that Eq. (118) has been used in Ref. [Sun2017] to compute the Coulomb matrix element $V_{PQ}(\mathbf{k})$ of Gaussians $P,Q$.
@@ Appendix H Spin-orbit coupling from HGH pseudopotentials
We employ spin-orbit coupling (SOC) from Hartwigsen-Goedecker-Hutter (HGH) pseudopotentials [Hartwigsen1998, Vogt2025],
@@
To demonstrate the effects of this correction, we provide in the following the PBE band structures of all monolayers with and without SOC (Fig. 10). We observe that SOC splits the top valence bands, predominantly at the K-point, in agreement with the literature [Gjerding2021].
@@ Appendix I GW band gaps with plasmon-pole model (BerkeleyGW)
In this Appendix, we present additional data of the band gaps for the TMDCs benchmarked in this work using BerkeleyGW (Table 2). We compare the generalized plasmon pole model (GPP) [Hybertsen1986] with the full frequency implementation (data of the full-frequency calculation is reproduced from Table 1). Some available results in the literature are given in Table 4.
@@
C2DB [Haastrup2018, Gjerding2021] GPAW PAW 18.1 3.184 12x12x1 N/A 59 $\infty$111extrapolated from 11, 13 and 15 Ry 2.53
Pela et al. [RodriguesPela2024] exciting LAPW+lo 16.0 3.190 18x18x1 400 N/A N/A 2.45
Echeverry et al. [Echeverry2016] VASP PAW 17.0 3.220 12x12x1 600 N/A N/A 2.31
Qiu et al. [Qiu2016] Berkeley GW NC 25.0 3.150 24x24x1 6000 350 35 2.45
@@
This work CP2K NC (TZV2P) 15.0 3.320 32x32x1 1.99
This work CP2K NC (aug-DZVP) 15.0 3.320 32x32x1 1.94
@@
C2DB [Haastrup2018, Gjerding2021] GPAW PAW 18.3 3.320 9x9x1 N/A 59 $\infty$1 2.12
Echeverry et al. [Echeverry2016] VASP PAW 17.0 3.320 12x12x1 600 N/A N/A 2.13
@@
This work CP2K NC (TZV2P) 15.0 3.186 32x32x1 2.44
This work CP2K NC (aug-DZVP) 15.0 3.186 32x32x1 2.34
@@
C2DB [Haastrup2018, Gjerding2021] GPAW PAW 18.1 3.186 12x12x1 N/A 59 $\infty$1 2.53
@@
This work CP2K NC (TZV2P) 15.0 3.319 32x32x1 2.02
This work CP2K NC (aug-DZVP) 15.0 3.319 32x32x1 1.93
@@
C2DB [Haastrup2018, Gjerding2021] GPAW PAW 18.3 3.319 12x12x1 N/A 59 $\infty$1 2.13
Echeverry et al. [Echeverry2016] VASP PAW 17.0 3.280 12x12x1 600 N/A N/A 2.06
@@
Gillen et al. [Roland2017] Berkeley GW NC 25.0 3.150 24x24x1 200 88 22 2.69
Kim et al. [Kim2021] Berkeley GW NC 25.0 3.190 12x12x1 3000 125 35 2.40
Qiu et al. [Qiu2016] Berkeley GW NC 25.0 3.150 24x24x1 6000 350 35 2.59
@@
Gillen et al. [Roland2017] Berkeley GW NC 25.0 3.279 24x24x1 200 88 22 2.33
Kim et al. [Kim2021] Berkeley GW NC 25.0 3.316 12x12x1 3000 125 35 2.08
@@
Figure 11: $G_{0}W_{0}$@PBE+SOC band gap of monolayer MoS2 and execution time as a function of the number of time points $\tau$ (Sec. IV), the $k$-mesh (Eqs. (21), (23), (32)), the filter threshold (for Eqs. (42), (56) and (59), see Eqs. (88) and (89)), the simulation cell box height (Sec. V.3), the size factor $\Delta$ for $V_{PQ}(\mathbf{k})$ (Eq. (110)) and the number of basis functions (Sec. V.3). Default parameters are reported on top. In (a), we also show $G_{0}W_{0}$@PBE+SOC with Hedin’s shift [hedin1965new, Hedin1999, Pollehn1998, Martin_Reining_Ceperley_2016, Golze2019, Golze2022] to avoid poles of the self-energy close to the quasiparticle solution [Veril2018, Schambeck2024].